### v1

@@ 3.1 Stimulus (IV) & Apparatus
To induce a cybersickness stimulus, we utilized an open-access simulator developed by [82]. This simulator immerses users in the perspective of a seated passenger in a cart on an amusement ride.
It is designed to provide a more controlled stimulus than typical virtual roller coasters (cf. [3]), while still maintaining a realistic and fun experience, as opposed to highly abstract approaches such as an optoelectronic drum (cf. [8]).
@@ 3.2 Procedure
Upon arrival at the first session, participants received both written and verbal explanations of the experiment. They then signed an informed consent form before being randomly assigned to either the control condition, which did not include a cybersickness stimulus, or the study condition, in which the stimulus was presented (see Section 3.1). A schematic representation of the experimental procedure is shown in Fig 2.
@@ 3.2.2 Control Condition
The control condition is particularly important for the endocrinological markers, as these are susceptible to environmental influences. Moreover, it became evident that subjective measurements such as the SSQ do not fully adhere to the zero-baseline assumption [7]. As a result, these measurements also become more robust when an appropriate control condition is implemented. We opted for a low-level (passive) control condition rather than closely mimicking the experimental condition. This decision was made to control for the well-known circadian rhythm of the HPA axis [35]. Participants in the control group were immediately directed to the resting room, where saliva samples and SSQ assessments were taken at the same time points as in the experimental condition. The first sample was labeled , following the notation of the study condition. They also completed the n-Back task at .
@@ 3.3.1 Questionnaires
The participants filled out the Simulator Sickness Questionnaire [28] (SSQ) at six time points, once before the experiment () and five times following the experiment (, , , , and ). All timestamps refer to the cortisol measurement, while the SSQ assessment always took place immediately before it, with a time offset of approximately 2 minutes. This instrument measures the symptoms related to cybersickness of nausea, disorientation, and oculomotor problems on a 4-point Likert scale from None to Severe using 16 items. It provides the three subscales of nausea, disorientation, and oculomotor as well as an overall score of cybersickness.
@@
Further, the participants were asked to rate their feelings of comfort throughout the study condition on a scale from 0 (not sick at all) to 10 (frank sickness). This scale is an adapted version of the Fast Motion Sickness Questionnaire [30] (FMS). It was adapted from 0 to 20 to make it more intuitive for the participants under the potentially stressful circumstances this experiment evokes (cf. [83, 1]). The FMS was used during the experiment to evaluate the current well-being of the participants during the stimulation. For this purpose, participants were asked to rate their current sickness every 30 seconds during the motion stimulation. Once they reached the threshold of 7, the stimulation was terminated to prevent the participants from severe physical reactions such as throwing up. One minute after the experiment, as well as with every SSQ following (study condition only), the participants gave another rating on the FMS.
@@
Furthermore, the participants were asked to fill out the Visually Induced Motion Sickness Susceptibility Questionnaire [32] (VIMSSQ) in the control condition between and . By using this instrument, we expect to evaluate the susceptibility of the participants to show symptoms related to motion sickness while using visual displays. It is expected to predict the SSQ values that the participants provide following the cybersickness-inducing experiment.
@@ 3.3.2 Salivary Cortisol and Alpha-Amylase
For this experiment, we used Salivettes (Sarstedt, Nümbrecht, Germany). Salivary cortisol levels were measured using a time-resolved fluorescence immunoassay [13]. Additionally, alpha-amylase activity was determined using 2-chloro-4-nitrophenyl-D-maltotriose (CNP-G3) as the relevant substrate.
The enzymatic action of alpha-amylase on this substrate yields 2-chloro-p-nitrophenol, which can be spectrophotometrically measured at 405 nm. The amount of alpha-amylase activity present in the sample is directly proportional to the increase in absorbance at 405 nm [44, 80]. The intra-assay Coefficient of Variation (CV) was 4.0-6.7%, the corresponding inter-assay CV was 7.1-9.0% for cortisol 2.8 - 6.3% (intra), and 5.5 – 7.6% (inter) for sAA. Analyses were performed in duplicate and averaged.
@@ 3.4 Participants
According to [34], we expected a medium effect size of in the psycho-endocrinological responses. To calculate the required sample size, we used G*power 3.1 [15], which resulted in a minimum of participants (, ). By listing this experiment on our intern study list as well as advertising through the E-Mail newsletter of the university, we were able to recruit the necessary amount of participants. Participants who smoked excessively ( cigarettes a day) (cf. [36, 5]), were pregnant, or took cortisol-containing medication were excluded from the study. Thirty-five healthy participants were enrolled in the study.
@@ 4.2 Salivary Cortisol and Alpha-Amylase
The repeated measures ANOVA revealed a significant main effect of condition, indicating that motion sickness had a statistically significant impact on salivary cortisol levels. Specifically, the analysis showed that participants exhibited significantly higher cortisol levels in the study condition compared to the control condition (, , ). A significant main effect of the sample was observed, demonstrating that cortisol levels varied across different samples (, , ). Additionally, the interaction between condition and sample was statistically significant (, , ), indicating that the effect of motion sickness on cortisol levels differed depending on the sample. These findings suggest that exposure to motion sickness induces a pronounced HPA axis response, as reflected in elevated cortisol levels (see Fig. 4a). In the study condition, the average response magnitude was (). Seventy percent of the participants showed an increase of at least and could thus be classified as responders (cf. [49, 37]).
@@ 5.1 SSQ Progression and Implications
First, we must acknowledge that the instrument used—the so-called Cybersicker—effectively induced cybersickness as intended. We observed a solid response compared to the baseline in both subjective measures, such as the SSQ and FMS, and objective measures, including cortisol, alpha-amylase, and, ultimately, the n-Back task. Fifteen minutes after the onset of stimulation, the average SSQ score in the study condition reached (), indicating a significantly elevated level (compared to , in the control condition). As expected, the results lie within a moderate range, also when compared to other studies (see [65] for a systematic overview). For direct comparison, [34] are cited again due to the close scientific alignment, reporting a total SSQ score of () pre-stimulation and () post-stimulation; note that no control condition was included in their study. This aligns with our design goal, which aimed to induce a level of cybersickness that still allowed participants to be systematically observed and interviewed after exposure. It is plausible that the immediate post-exposure response was even higher. However, we deliberately chose to standardize the measurement time across all participants relative to the onset of the stimulus rather than its conclusion. Together with the necessary follow-up for biophysical measurements and VR offboarding, this resulted in a delay of the first SSQ measurement post-stimulation by approximately 7 to 12 minutes, as previously mentioned—depending on whether and when the session was terminated. The shortest exposure was recorded for a subject who reached an FMS score of 7 in just under 3 minutes. Therefore, any statements regarding the period immediately following the exposure and the first measurement remain speculative.
When examining the SSQ subscales, the strongest response is found in the Nausea and Disorientation scales, consistent with other studies using exposure durations comparable to the one applied here [59].
@@ 5.1.1 Delayed Symptoms
One of the motivations for this study and the post-hoc analysis of cybersickness exposure was the recurring anecdotal reports of delayed symptoms—such as headaches—that appear in some individuals with a considerable time lag after exposure [81]. This phenomenon is particularly relevant when considering safety aspects of VR usage (see Section 5.3), but it also raises important methodological concerns. Delayed symptom onset could compromise the validity of within-subject study designs by introducing uncontrolled, higher-order carry-over effects [85].
@@
It should also be noted that, due to the study design, we cannot rule out the possibility that the n-Back task administered after the 30-minute SSQ measurement may have interacted with or influenced the reported symptoms at 45, 60, and 90 minutes (cf. [62]). Replication of the results under controlled variation of this factor is therefore necessary.
@@ 5.1.2 Time of Recovery
Another notable observation is that, on average, participants returned to the SSQ baseline level of the control condition after 90 minutes. This is a considerable duration, aligning with the few long-term observations reported in related studies (see Section 2.2) and raising concerns about washout periods in within-subjects experiments, which are typically much shorter. However, it is important to note that the n-Back task, as just mentioned, conducted at minutes after stimulation onset (approximately 30 minutes post-stimulation), may have interacted with participants’ SSQ score through items such as fullness of head. On the other hand, the fact that our stimulation protocol resulted in 70% nausea-induced early terminations suggests that recovery times could be even longer under stronger stimulation conditions.
@@ 5.2 Cortisol Reaction - Cybersicker as a Stressor?
Based on findings from previous studies, we expected a stress response including self-reported stress, elevated salivary cortisol as well as autonomic activation [34]. Overall, we found the expected pattern of increased subjective stress and physiological responses. However, given that cortisol is primarily considered a marker for social stress [12], the average magnitude of the baseline-to-peak response—approximately 300% of baseline—is surprisingly high and has rarely been found in stress induction studies, neither with an in-person settings [78], nor in VR setups [86, 43]. However, it should be taken into account that the individual responses in the present study vary significantly. Taking this high variability into account, we found a baseline-peak effect size of Cohen’s , which is roughly comparable to previous studies with psychological stressors (such as the TSST) and previous studies inducing cybersickness.
For instance, [34] observed a cortisol level before cybersickness induction of () and after stimulation of (), corresponding to an average increase of 37%. Due to the lower variability, this corresponds to an effect size of Cohen’s . In comparison, meta-analyses summarizing TSST studies reported overall pre-post effect sizes of approximately Hedges - [20, 61], while studies using VR versions of the TSST observed slightly lower effect sizes around [22].
@@
Furthermore, we observed a delayed response of salivary cortisol compared to the SSQ, with peaks occurring at and after the start of the simulation. This is expected, as cortisol is released with a temporal delay relative to the stressor and can be measured in saliva with an additional delay [20]. However, 90 minutes after the experiment began, cortisol levels had still not returned to control levels and remained almost three times higher than baseline. This extended recovery period may be due to the strong stress response (with a pronounced cortisol increase) in our study or related to other aspects of the Cybersicker like physiological stress responses in other systems, e.g., the cardiovascular system [34, 2].
@@
In the present study, we observed that regardless of the cognitive load of the working memory task, the induction of cybersickness led to a significant drop in performance compared to the control condition. One explanation for the reduced working memory performance after cybersickness induction relates to the well-established effects of delayed effects of stress-induced cortisol secretion on working memory [19, 18]. Presumably, elevated cortisol levels approximately 30 minutes after stress induction might have impaired working memory performance, as reflected in reduced accuracy on the n-Back task compared to the control condition. This finding aligns with previous research demonstrating that acute stress and the subsequent glucocorticoid response can transiently disrupt prefrontal cortex functioning, which is critical for working memory processes [64]. Cortisol’s delayed effects may interfere with neural activity in the dorsolateral prefrontal cortex, leading to diminished cognitive flexibility and attentional control [4].

### v2

@@ 3.1 Stimulus (IV) & Apparatus
To induce a cybersickness stimulus, we utilized an open-access simulator developed by [zielasko2024cybersicker]. This simulator immerses users in the perspective of a seated passenger in a cart on an amusement ride.
It is designed to provide a more controlled stimulus than typical virtual roller coasters (cf. [Ang2023]), while still maintaining a realistic and fun experience, as opposed to highly abstract approaches such as an optoelectronic drum (cf. [Bubka2006Optokinetic]).
@@ 3.2 Procedure
Upon arrival at the first session, participants received both written and verbal explanations of the experiment. They then signed an informed consent form before being randomly assigned to a sequence-balanced order of the two experimental sessions: a control condition without a cybersickness stimulus and a study condition in which the stimulus was presented (see Section 3.1). A schematic representation of the experimental procedure is shown in Fig. 2.
@@ 3.2.2 Control Condition
The control condition is particularly important for the endocrinological markers, as these are susceptible to environmental influences. Moreover, it became evident that subjective measurements such as the SSQ do not fully adhere to the zero-baseline assumption [Brown2022zeroBaseline]. As a result, these measurements also become more robust when an appropriate control condition is implemented. We deliberately chose a low-level (passive), time-matched control condition because our primary outcomes included endocrine markers related to HPA-axis activity, which are highly sensitive to contextual arousal, novelty, expectancy, and environmental stimulation. Any VR exposure — even a “less sickening” or stationary version — would still constitute an active manipulation (e.g., visual stimulation, immersion, novelty, equipment setup), potentially influencing HPA-axis responses independent of cybersickness. The passive control was therefore intended to provide a conservative baseline while controlling for the well-known circadian rhythm of the HPA axis [Kirschbaum1994] as well as repeated sampling effects. Accordingly, participants in the control session followed the identical measurement schedule (saliva sampling and SSQ assessments at the same time points as in the study session) while resting in the recovery room. During the control session, participants were immediately directed to the resting room, where saliva samples and SSQ assessments were taken at the same time points as in the study session. The first sample was labeled , following the notation of the study condition. They also completed the n-Back task at .
@@ 3.3.1 Questionnaires
The participants filled out the Simulator Sickness Questionnaire (SSQ) at six time points, once before the experiment () and five times following the experiment (, , , , and ). All timestamps refer to the cortisol measurement, while the SSQ assessment always took place immediately before it, with a time offset of approximately 2 minutes. This instrument measures the symptoms related to cybersickness of nausea, disorientation, and oculomotor problems on a 4-point Likert scale from None to Severe using 16 items. It provides the three subscales of nausea, disorientation, and oculomotor as well as an overall score of cybersickness.
@@
Further, the participants were asked to rate their feelings of comfort throughout the study condition on a scale from 0 (not sick at all) to 10 (frank sickness). This scale is an adapted version of the Fast Motion Sickness Questionnaire (FMS). It was adapted from 0 to 20 to make it more intuitive for the participants under the potentially stressful circumstances this experiment evokes (cf. [zielasko2016hmdnav, Adhikari2022]). The FMS was used during the experiment to evaluate the current well-being of the participants during the stimulation. For this purpose, participants were asked to rate their current sickness every 30 seconds during the motion stimulation. Once they reached the threshold of 7, the stimulation was terminated to prevent the participants from severe physical reactions such as throwing up. One minute after the experiment, as well as with every SSQ following (study condition only), the participants gave another rating on the FMS.
@@
Furthermore, the participants were asked to fill out the Visually Induced Motion Sickness Susceptibility Questionnaire (VIMSSQ) [Keshavarz2023vimssq] in the control condition between and . By using this instrument, we expect to evaluate the susceptibility of the participants to show symptoms related to motion sickness while using visual displays. It is expected to predict the SSQ values that the participants provide following the cybersickness-inducing experiment.
@@ 3.3.2 Salivary Cortisol and Alpha-Amylase
For this experiment, we used Salivettes (Sarstedt, Nümbrecht, Germany). Salivary cortisol levels were measured using a time-resolved fluorescence immunoassay [Dressendorfer1992CortisolBiotin]. Additionally, alpha-amylase activity was determined using 2-chloro-4-nitrophenyl-D-maltotriose (CNP-G3) as the relevant substrate.
The enzymatic action of alpha-amylase on this substrate yields 2-chloro-p-nitrophenol, which can be spectrophotometrically measured at 405 nm. The amount of alpha-amylase activity present in the sample is directly proportional to the increase in absorbance at 405 nm [LorentzGutschowRenner1999, WinnDeen1988]. The intra-assay Coefficient of Variation (CV) was 4.0-6.7%, the corresponding inter-assay CV was 7.1-9.0% for cortisol 2.8 - 6.3% (intra), and 5.5 – 7.6% (inter) for sAA. Analyses were performed in duplicate and averaged.
@@ 3.4 Participants
Given the limited and heterogeneous evidence on psycho-endocrinological and psychophysiological correlates of cybersickness, this study was designed as largely exploratory. The sample size was therefore determined based on prior related work rather than a single confirmatory hypothesis. Following [Kim2022], we used G*Power 3.1 [Faul2007gPower] to obtain an initial orientation for the required sample size, assuming a medium effect size (). This resulted in a minimum sample size of participants (, ). By listing this experiment on our intern study list as well as advertising through the E-Mail newsletter of the university, we were able to recruit the necessary amount of participants. Participants who smoked excessively ( cigarettes a day) (cf. [Kirschbaum1992, Badrick2007]), were pregnant, or took cortisol-containing medication were excluded from the study. In total, thirty-five healthy participants were enrolled to account for potential dropouts and to increase robustness given individual differences in cybersickness susceptibility.
@@
Susceptibility to visually induced motion sickness was assessed using the VIMSSQ. The sample showed substantial interindividual variability (VIMSSQ total score: , , range: 1.1–103.9), indicating a heterogeneous group spanning low to highly susceptible individuals rather than a sample skewed toward extreme susceptibility.
@@ 4.2 Salivary Cortisol and Alpha-Amylase
The repeated measures ANOVA revealed a significant main effect of condition, indicating that motion sickness had a statistically significant impact on salivary cortisol levels. Specifically, the analysis showed that participants exhibited significantly higher cortisol levels in the study condition compared to the control condition (, , ). A significant main effect of the sample was observed, demonstrating that cortisol levels varied across different samples (, , ). Additionally, the interaction between condition and sample was statistically significant (, , ), indicating that the effect of motion sickness on cortisol levels differed depending on the sample. These findings suggest that exposure to motion sickness induces a pronounced HPA axis response, as reflected in elevated cortisol levels (see Fig. 4a). In the study condition, the average response magnitude was (). Seventy percent of the participants showed an increase of at least and could thus be classified as responders (cf. [miller2013classification, kudielka2007ten]).
@@ 5.1 SSQ Progression and Implications
First, we must acknowledge that the instrument used—the so-called Cybersicker—effectively induced cybersickness as intended. We observed a solid response compared to the baseline in both subjective measures, such as the SSQ and FMS, and objective measures, including cortisol, alpha-amylase, and, ultimately, the n-Back task. Fifteen minutes after the onset of stimulation, the average SSQ score in the study condition reached (), indicating a significantly elevated level (compared to , in the control condition). As expected, the results lie within a moderate range, also when compared to other studies (see [SimnnVicente2024] for a systematic overview). For direct comparison, [Kim2022] are cited again due to the close scientific alignment, reporting a total SSQ score of () pre-stimulation and () post-stimulation; note that no control condition was included in their study. This aligns with our design goal, which aimed to induce a level of cybersickness that still allowed participants to be systematically observed and interviewed after exposure. It is plausible that the immediate post-exposure response was even higher. However, we deliberately chose to standardize the measurement time across all participants relative to the onset of the stimulus rather than its conclusion. Together with the necessary follow-up for biophysical measurements and VR offboarding, this resulted in a delay of the first SSQ measurement post-stimulation by approximately 7 to 12 minutes, as previously mentioned—depending on whether and when the session was terminated. The shortest exposure was recorded for a subject who reached an FMS score of 7 in just under 3 minutes. Therefore, any statements regarding the period immediately following the exposure and the first measurement remain speculative.
When examining the SSQ subscales, the strongest response is found in the Nausea and Disorientation scales, consistent with other studies using exposure durations comparable to the one applied here [Saredakis2020factors].
@@ 5.1.1 Delayed Symptoms
One of the motivations for this study and the post-hoc analysis of cybersickness exposure was the recurring anecdotal reports of delayed symptoms—such as headaches—that appear in some individuals with a considerable time lag after exposure [Zielasko2021sickness]. This phenomenon is particularly relevant when considering safety aspects of VR usage (see Section 5.3), but it also raises important methodological concerns. Delayed symptom onset could compromise the validity of within-subject study designs by introducing uncontrolled, higher-order carry-over effects [zielasko2024carryOver].
@@
The broad distribution of VIMSSQ scores in the present sample further supports the interpretation that these delayed effects are not driven by a narrowly selected high-susceptibility subgroup, but rather reflect genuine interindividual variability in responses to the cybersickness stimulus.
It should also be noted that, due to the study design, we cannot rule out the possibility that the n-Back task administered after the 30-minute SSQ measurement may have interacted with or influenced the reported symptoms at 45, 60, and 90 minutes (cf. [Sepich2022]). Replication of the results under controlled variation of this factor is therefore necessary.
@@ 5.1.2 Time of Recovery
Another notable observation is that, on average, participants returned to the SSQ baseline level of the control condition after 90 minutes. This is a considerable duration, aligning with the few long-term observations reported in related studies (see Section 2.2) and raising concerns about washout periods in within-subjects experiments, which are typically much shorter. However, it is important to note that the n-Back task, as just mentioned, conducted at minutes after stimulation onset (approximately 30 minutes post-stimulation), may have interacted with participants’ SSQ score through items such as fullness of head. On the other hand, the fact that our stimulation protocol resulted in 70% nausea-induced early terminations suggests that recovery times could be even longer under stronger stimulation conditions. An important aspect in interpreting post-exposure recovery is the effective dose of cybersickness experienced by individual participants. Cybersickness severity is known to depend on the characteristics and duration of the stimulus as well as on individual susceptibility. Based on established categorizations of SSQ scores [Stanney1997], the average sickness levels observed in the present study fall within the high to severe range. This provides a plausible explanation for the pronounced and prolonged recovery times observed in a subset of participants. At the same time, early termination of the stimulation by some participants resulted in a lower effective dose, which is consistent with reports of rapid recovery in other studies. Taken together, these findings support the notion that post-exposure recovery dynamics should be interpreted as dose-dependent rather than uniform across participants.
@@ 5.2 Cortisol Reaction - Cybersicker as a Stressor?
Based on findings from previous studies, we expected a stress response including self-reported stress, elevated salivary cortisol as well as autonomic activation [Kim2022]. Overall, we found the expected pattern of increased subjective stress and physiological responses. However, given that cortisol is primarily considered a marker for social stress [Dickerson2004acute], the average magnitude of the baseline-to-peak response—approximately 300% of baseline—is surprisingly high and has rarely been found in stress induction studies, neither with an in-person settings [vonDawans2021], nor in VR setups [Zimmer2019, linnig2025openTSST]. However, it should be taken into account that the individual responses in the present study vary significantly. Taking this high variability into account, we found a baseline-peak effect size of Cohen’s , which is roughly comparable to previous studies with psychological stressors (such as the TSST) and previous studies inducing cybersickness.
For instance, [Kim2022] observed a cortisol level before cybersickness induction of () and after stimulation of (), corresponding to an average increase of 37%. Due to the lower variability, this corresponds to an effect size of Cohen’s . In comparison, meta-analyses summarizing TSST studies reported overall pre-post effect sizes of approximately Hedges - [Goodman2017, Seel2025], while studies using VR versions of the TSST observed slightly lower effect sizes around [Helminen2019].
Importantly, our findings are in line with recent work directly comparing motion sickness and psychological stress in terms of cortisol responses. Golding and colleagues reported that motion sickness induced by cross-coupled Coriolis motion produced substantially larger increases in salivary cortisol than the Trier Social Stress Test (TSST), despite comparable levels of self-reported stress and anxiety between conditions [Golding2025cortisolMotion]. This dissociation suggests that motion sickness may act as a potent physiological stressor, engaging HPA-axis activity through mechanisms that are at least partly independent of purely psychological stress appraisal. The pronounced and prolonged cortisol elevations observed following VR-induced motion sickness in the present study are consistent with this interpretation and support the view that cybersickness-related stress responses may differ qualitatively from classic psychosocial stress paradigms.
@@
Furthermore, we observed a delayed response of salivary cortisol compared to the SSQ, with peaks occurring at and after the start of the simulation. This is expected, as cortisol is released with a temporal delay relative to the stressor and can be measured in saliva with an additional delay [Goodman2017]. However, 90 minutes after the experiment began, cortisol levels had still not returned to control levels and remained almost three times higher than baseline. This extended recovery period may be due to the strong stress response (with a pronounced cortisol increase) in our study or related to other aspects of the Cybersicker like physiological stress responses in other systems, e.g., the cardiovascular system [Kim2022, Allen2014].
@@
In the present study, we observed that regardless of the cognitive load of the working memory task, the induction of cybersickness led to a significant drop in performance compared to the control condition. One explanation for the reduced working memory performance after cybersickness induction relates to the well-established effects of delayed effects of stress-induced cortisol secretion on working memory [Geissler2023, Geissler2025]. Presumably, elevated cortisol levels approximately 30 minutes after stress induction might have impaired working memory performance, as reflected in reduced accuracy on the n-Back task compared to the control condition. This finding aligns with previous research demonstrating that acute stress and the subsequent glucocorticoid response can transiently disrupt prefrontal cortex functioning, which is critical for working memory processes [Shields2016]. Cortisol’s delayed effects may interfere with neural activity in the dorsolateral prefrontal cortex, leading to diminished cognitive flexibility and attentional control [Arnsten2009StressPrefrontal].
@@
The observed impairments in cognitive performance following cybersickness induction are consistent with findings from studies employing other forms of symptom modulation. Previous work has shown that both central stimulation (e.g., visually or vestibularly driven perturbations; cf. [Benelli2023cybersickness]) and peripheral stimulation can induce comparable effects on discomfort and cognitive performance. This suggests that the cognitive aftereffects observed in the present study may reflect a more general response to physiological and sensory stressors rather than a mechanism specific to cybersickness alone.